I had this question on a test and missed it.
It says;
Use the converse of the pythagorean Theorem to find the triples that could be the sides of a right triangle from the given choices.
for the number one the choices are
a.12,16,18
b.15,17,4
c.40,399,400
d.12,16,20
and for the second problem the are
a.40,75,85
b.60,80,90
c.13,12,12
d.51,49,7

Answer by stanbon(75887)

Use the converse of the pythagorean Theorem to find the triples that could be the sides of a right triangle from the given choices.
for the number one the choices are
a.12,16,18
b.15,17,4
c.40,399,400
d.12,16,20
---
Ans: d because 20^2 = 12^2 + 16^2
----------------------------------
and for the second problem the are
a.40,75,85
b.60,80,90
c.13,12,12
d.51,49,7
---
Ans: a because 85^2 = 75^2 + 40^2
======================================
Cheers,
Stan H.